OpenStudy (anonymous):

32/7 as a mixed number

OpenStudy (anonymous):

\[4\frac{ 4 }{ 7 }\]

OpenStudy (anonymous):

32/7 = 4 4/7 32/7 = 4(the whole number) remainder 4(your numerator) over your original denominator 7.

OpenStudy (anonymous):

first divide numerator by denominator and what comes as integer write in middle and place the divisor aside in left below the line and remainder above the line in this case when we divide 32 by 7 we get 4 as quotient and 4 as remainder and put it in the above described manner.making it 4 4/7